OpenStudy (anonymous):

find the rate of change of V if; V = x^2 y^3 x=cosbt and y=sinbt

OpenStudy (anonymous):

V=x^2*y^3 and x and y are functions of t. so now , dv/dt=2x*y^3*dx/dt+3y^2*x^2*dy/dt u know dx/dt=-bsin(bt) and dy/dt=bcos(bt) , substitue and get the answer. see i've done only chain differentiation. :)